OpenStudy (anonymous):

The areas of two similar rectangles are 80 square feet and 500 square feet. What scale factor applied to the smaller rectangle will give the larger rectangle? A.25/64 B.5/8 C.25/4 D.5/2

OpenStudy (anonymous):

I think it's D.....

OpenStudy (anonymous):

what do you think @mwill<3

OpenStudy (anonymous):

Your first instinct is usually correct.

OpenStudy (anonymous):

what do you think @timaashorty

OpenStudy (ipwnbunnies):

Make it into an equation. 80x = 500

OpenStudy (ipwnbunnies):

Solve for x, simplify the fraction. That'll be your scale factor.

OpenStudy (timaashorty):

I think it's C.

OpenStudy (timaashorty):

especially if you do what ipwn said then x = 25/4

OpenStudy (anonymous):

i thought I had to square root 25/4...

OpenStudy (timaashorty):

or find the ratios by simplifying : 500/80 = 50/8 = 25/4

OpenStudy (ipwnbunnies):

I don't think you have to because we're already finding the scale factor of the areas.

OpenStudy (anonymous):

oh thank you @timaashorty & @iPwnBunnies :)

OpenStudy (timaashorty):

You're very welcome (:

OpenStudy (ipwnbunnies):

Good observation though, if we were given side length and had to find scale factor of area, we might've used square root/squaring.

OpenStudy (anonymous):

thanks for the medal @iPwnBunnies

OpenStudy (anonymous):

the answer was D....

OpenStudy (anonymous):

What was the answer? I have 2/5 as an answer option, but not 5/2.